Assume that you have been given the following information on Purcell Industries' call options:

Current stock price = $15 Strike price of option = $14
Time to maturity of option = 9 months Risk-free rate = 8%
Variance of stock return = 0.13
d1 = 0.56923 N(d1) = 0.71540
d2 = 0.25698 N(d2) = 0.60140

According to the Black-Scholes option pricing model, what is the option's value? Do not round intermediate calculations. Round your answer to the nearest cent. Use only the values provided in the problem statement for your calculations.
Answer #1

Call Value =Stock Price* N(d1)–Strike ×e–Rt × N(d2) =15*0.71540-14*EXP(-8%*9/12)*0.60140 =2.80

Call Value =2.80